# An algebraic theory of Łojasiewicz exponents

### v1

@@ 4.2. Algebraic and analytic Łojasiewicz exponents
This is the Lejeune-Jalabert–Teissier analytic criterion for integral closure; see [36, Théorème 7.2]. For the reader’s convenience and to fix normalization conventions, we include a self-contained argument in the appendix. ∎
@@ 4.3. Exponents at infinity in the algebraic framework
Equivalently, $|u_{Y}|^{q}\leq C^{q}|u_{X}|^{p}$ holds in a punctured neighborhood of $\xi$. By the Lejeune-Jalabert–Teissier criterion, Lemma 4.4, this is equivalent to the containment
$(u_{Y})^{q}\subseteq\overline{(u_{X})^{p}}\text{ in }\mathcal{O}_{\overline{\Gamma}_{F},\xi}.$
Since $D_{X}$ and $D_{Y}$ are Cartier divisors, we have $\mathscr{I}_{X,\xi}=(u_{X})$ and $\mathscr{I}_{Y,\xi}=(u_{Y})$. Thus, the above containment is equivalent to
@@ 5.2. Valuative criterion for integral closure
This result is classical and goes back to Rees [45] and Lipman [38]; see also Huneke–Swanson [30, Chapters 6 and 10] for a modern account. ∎
@@
with $A,B$ effective Cartier divisors supported on the exceptional locus.
@@
In particular, integral–closure containments are completely determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E}\}$ arising from the exceptional divisors of $\pi$.
@@
Observe that $\mathfrak{a}^{p}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-pA)$ and $\mathfrak{b}^{q}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-qB)$ are invertible, and so, integrally closed on $Y$. Furthermore, since $Y$ is normal, $\pi$ factors through the normalized blowup of $\mathfrak{a}$. Thus, we have the following standard equality (cf. [30, Chapters 10 and 19]):
@@ 5.4. From supremum to maximum: finite testing and Noetherian Rees algebra
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of ideals in $R$. Assume that the Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}}):=\bigoplus_{p\geq 0}\mathfrak{a}_{p}T^{p}$ is finitely generated as an RR–algebra. Let $\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}$ denote its integral closure, and set $\overline{Y}:=\operatorname{Proj}\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}.$ Then, there exists a finite set of divisorial valuations
@@
arising from prime divisors on $\overline{Y}$ dominating $\operatorname{Spec}R$, such that for every filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in $R$ satisfying $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $1\leq i\leq r$, one has
@@
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration whose Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated. Then, there exists a finite set of divisorial valuations $\mathcal{V}$ with the following property:
@@
By Theorem 5.9, there exists a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})$ computing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ for all filtrations $\mathfrak{b}_{\bullet}$}. The valuative criterion for integral closure then implies that the same set of valuations detects all containments $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$, proving the claim. ∎
@@ 5.5. From supremum to maximum: normalized valuation space and linearly boundedness
By the discussion following [9, Definition 1.3] (see also [31, Proposition 5.9] and the discussion that followed), the normalized space $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ is compact in the topology of pointwise convergence on $R$.
Note that, by [31, Lemma 4.1], $v\mapsto v(\mathfrak{m})$ and $v\mapsto v(\mathfrak{q})$ are continuous on $\operatorname{Val}_{\mathfrak{m}}$. Define
@@
We work on the normalized slice $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})=\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$, so $v(\mathfrak{m})=1$ for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. We shall also use notations and terminology from Favre–Jonsson [16] monograph.
Step 1: a non-divisorial end and divisorial approximants. By Favre–Jonsson’s classification of the valuative tree of $R=\mathbb{C}[[x,y]]$ in [16], there exist non-divisorial ends (for instance curve valuations or infinitely singular valuations). Fix such an end valuation $\nu\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Then, $\nu$ is not divisorial, and there exists an increasing sequence of divisorial valuations
$\nu_{1}\prec\nu_{2}\prec\cdots\prec\nu_{i}\prec\cdots\prec\nu$
converging to $\nu$ in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, where $\prec$ denotes the tree order. We use the fact that the skewness function $\alpha(\cdot)$ on the valuative tree is strictly increasing along any nontrivial segment (see [16]).
Step 2: realizing $\alpha(v\wedge\nu_{i})$ as $v(J_{i})$. For valuations $v,\mu\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, Favre–Jonsson [16] define the intersection product by
$v\cdot\mu:=\alpha(v\wedge\mu),$
and for fixed $\mu$ the associated extremal tree potential is $g_{\mu}(v):=v\cdot\mu=\alpha(v\wedge\mu)$. If $\mu$ is divisorial, then it was proved (see the proof of [17, Prop. 5.15]; although this was written for the ring $\mathbb{C}\{x,y\}$ of convergent formal power series, the argument applies verbatim to the formal power series ring $\mathbb{C}[[x,y]]$) that there exist irreducible elements $\phi_{\mu,1},\phi_{\mu,2}\in\mathfrak{m}$ such that
$g_{\mu}(v)=\min\{v(\phi_{\mu,1}),v(\phi_{\mu,2})\}\quad\text{for all }v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}).$
Applying this to $\mu=\nu_{i}$, we may choose irreducibles $\phi_{i,1},\phi_{i,2}\in\mathfrak{m}$ with
$\alpha(v\wedge\nu_{i})=\min\{v(\phi_{i,1}),v(\phi_{i,2})\}\quad\forall\,v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}).$
Set
$J_{i}:=(\phi_{i,1},\phi_{i,2})\subseteq R.$
Since $\phi_{i,1},\phi_{i,2}$ are distinct irreducible germs through the origin, the ideal $J_{i}$ has height $2$ in the regular local ring $R=\mathbb{C}[[x,y]]$ and is therefore $\mathfrak{m}$–primary. For all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ we have the exact identity
(5.4) $v(J_{i})=\alpha(v\wedge\nu_{i}).$
Set $A_{i}:=\alpha(\nu_{i})\in(0,\infty)$.
Step 3: normalized coefficients and definition of the filtrations. Set
$c_{i}:=\frac{2^{-i}}{A_{i}}.$
Since $\alpha(\operatorname{\mathrm{ord}}_{\mathfrak{m}})=1$ on the normalized slice and $\alpha$ is increasing, we have $A_{i}\geq 1$ and so $c_{i}\leq 2^{-i}$. Using (5.4), for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$,
$0\leq c_{i}\,v(J_{i})=c_{i}\,\alpha(v\wedge\nu_{i})\leq c_{i}\,\alpha(\nu_{i})=2^{-i}.$
Define ${\mathfrak{b}_{\bullet}}:=\mathfrak{m}^{\bullet}$, i.e., $\mathfrak{b}_{p}=\mathfrak{m}^{p}$. Define ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ by
(5.5) $\mathfrak{a}_{p}:=\mathfrak{m}^{p}\cdot\prod_{i\geq 1}J_{i}^{\lfloor p\,c_{i}\rfloor}.$
Since $c_{i}\leq 2^{-i}$, we have $\lfloor p\,c_{i}\rfloor=0$ for all $i\gg 1$, so the product is finite for each $p$. In particular, each $\mathfrak{a}_{p}$ is $\mathfrak{m}$–primary. Observe further that, for all $p,p^{\prime}\geq 1$ and all $i$, $\lfloor(p+p^{\prime})c_{i}\rfloor\geq\lfloor pc_{i}\rfloor+\lfloor p^{\prime}c_{i}\rfloor.$ Therefore,
$\mathfrak{a}_{p}\mathfrak{a}_{p^{\prime}}=\mathfrak{m}^{p+p^{\prime}}\cdot\prod_{i}J_{i}^{\lfloor pc_{i}\rfloor+\lfloor p^{\prime}c_{i}\rfloor}\subseteq\mathfrak{m}^{p+p^{\prime}}\cdot\prod_{i}J_{i}^{\lfloor(p+p^{\prime})c_{i}\rfloor}=\mathfrak{a}_{p+p^{\prime}}.$
Moreover, $\mathfrak{a}_{p+1}\subseteq\mathfrak{a}_{p}$, so $\mathfrak{a}_{\bullet}$} is a filtration of ideals.
Step 4: computation of $v({\mathfrak{a}_{\bullet}})$. Fix $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. By definition, we have
$v(\mathfrak{a}_{p})=v(\mathfrak{m}^{p})+\sum_{i\geq 1}\lfloor p\,c_{i}\rfloor\,v(J_{i})=p\,v(\mathfrak{m})+\sum_{i\geq 1}\lfloor p\,c_{i}\rfloor\,v(J_{i}).$
Dividing by $p$,
$\frac{v(\mathfrak{a}_{p})}{p}=v(\mathfrak{m})+\sum_{i\geq 1}\frac{\lfloor p\,c_{i}\rfloor}{p}\,v(J_{i}).$
For each $i$, $\frac{\lfloor p\,c_{i}\rfloor}{p}\to c_{i}$ as $p\to\infty$, and
$0\leq\frac{\lfloor p\,c_{i}\rfloor}{p}\,v(J_{i})\leq c_{i}\,v(J_{i})\leq 2^{-i}.$
Since $\sum_{i\geq 1}2^{-i}<\infty$, dominated convergence applies and yields
$v({\mathfrak{a}_{\bullet}})=\lim_{p\to\infty}\frac{v(\mathfrak{a}_{p})}{p}=v(\mathfrak{m})+\sum_{i\geq 1}c_{i}\,v(J_{i})=v(\mathfrak{m})+\sum_{i\geq 1}c_{i}\,\alpha(v\wedge\nu_{i}).$
Since ${\mathfrak{b}_{\bullet}}=\mathfrak{m}^{\bullet}$ and $v(\mathfrak{m})=1$ on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, we have $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})=v({\mathfrak{a}_{\bullet}})$ on this slice.
Step 5: unique maximizer and non-attainment by divisorial valuations. Because $\nu\wedge\nu_{i}=\nu_{i}$ for all $i$,
$\nu({\mathfrak{a}_{\bullet}})=1+\sum_{i\geq 1}c_{i}\,\alpha(\nu_{i})=1+\sum_{i\geq 1}2^{-i}=2.$
Let $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ with $v\neq\nu$. Since $\nu$ is an end and $\nu_{i}\uparrow\nu$, there exists $i_{0}$ such that $v\not\succeq\nu_{i_{0}}$; otherwise $v\succeq\nu_{i}$ for all $i$ and so $v\succeq\nu$, forcing $v=\nu$. For this $i_{0}$ we have $v\wedge\nu_{i_{0}}\prec\nu_{i_{0}}$, and strict monotonicity of $\alpha$ along segments implies
$\alpha(v\wedge\nu_{i_{0}})<\alpha(\nu_{i_{0}}).$
For all $i$, $\alpha(v\wedge\nu_{i})\leq\alpha(\nu_{i})$, with strict inequality at $i=i_{0}$. Therefore,
$v({\mathfrak{a}_{\bullet}})=1+\sum_{i\geq 1}c_{i}\,\alpha(v\wedge\nu_{i})<1+\sum_{i\geq 1}c_{i}\,\alpha(\nu_{i})=2=\nu({\mathfrak{a}_{\bullet}}).$
Thus, $\nu$ is the unique maximizer of $v({\mathfrak{a}_{\bullet}})$ on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ and, hence, of $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})$. Since $\nu$ is not divisorial, no divisorial valuation attains the supremum.
@@
Step 6: linear boundedness. Since ${\mathfrak{b}_{\bullet}}=\mathfrak{m}^{\bullet}$, it is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$. For $\mathfrak{a}_{\bullet}$}, the definition (5.5) gives $\mathfrak{m}^{p}\subseteq\mathfrak{a}_{p}$ for all $p$. Using $J_{i}\subseteq\mathfrak{m}$ and $\sum_{i}c_{i}\leq 1$, we obtain
$\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}\cdot\prod_{i}\mathfrak{m}^{\lfloor p\,c_{i}\rfloor}=\mathfrak{m}^{p+\sum_{i}\lfloor p\,c_{i}\rfloor}\subseteq\mathfrak{m}^{p+\lfloor p\sum_{i}c_{i}\rfloor}\subseteq\mathfrak{m}^{2p}.$
Hence, $\mathfrak{a}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$, and the hypotheses of Theorem 5.14 are satisfied. This completes the proof. ∎

### v2

@@ 1. Introduction
Our approach also resonates with the valuation-theoretic framework for multiplier ideals, plurisubharmonic singularities, and interpolation developed by Boucksom–Favre–Jonsson [12], Favre–Jonsson [19, 20, 21], and more recently by Bao–Guan–Mi–Yuan [1, 2, 3]. While our setting is algebraic and centers on ideal containment, a common underlying theme is that singularity exponents can be described as optimization problems over suitable spaces of valuations. In this sense, the present algebraic theory of Łojasiewicz exponents runs parallel to the valuative descriptions of multiplier ideals, analytic singularities, and interpolation phenomena studied in the above works.
@@ 4.2. Algebraic and analytic Łojasiewicz exponents
This is the Lejeune-Jalabert–Teissier analytic criterion for integral closure; see [40, Théorème 7.2]. For the reader’s convenience and to fix normalization conventions, we include a self-contained argument in the appendix. ∎
@@ 4.3. Exponents at infinity in the algebraic framework
Equivalently, $|u_{Y}|^{q}\leq C^{q}|u_{X}|^{p}$ holds in a punctured neighborhood of $\xi$. By [40, Théorèm 2.1], this is the case if and only if $(u_{Y})^{q}\subseteq\overline{(u_{X})^{p}}$. Since $D_{X}$ and $D_{Y}$ are Cartier divisors, we have $\mathscr{I}_{X,\xi}=(u_{X})$ and $\mathscr{I}_{Y,\xi}=(u_{Y})$. Thus, the above containment is equivalent to
@@ 5.2. Valuative criterion for integral closure
This result is classical and goes back to Rees [49] and Lipman [42]; see also Huneke–Swanson [34, Chapters 6 and 10] for a modern account. ∎
@@
with $A,B$ effective Cartier divisors on $Y$.
@@
In particular, integral–closure containments are completely determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E}\}$ arising from prime divisors in the supports of $A$ and $B$.
@@
Observe that $\mathfrak{a}^{p}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-pA)$ and $\mathfrak{b}^{q}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-qB)$ are invertible, and so, integrally closed on $Y$. Furthermore, since $Y$ is normal, $\pi$ factors through the normalized blowup of $\mathfrak{a}$. Thus, we have the following standard equality (cf. [34, Chapters 10 and 19]):
@@ 5.4. From supremum to maximum: finite testing and Noetherian Rees algebra
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of $\mathfrak{m}$-primary ideals in $R$. Assume that the Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}}):=\bigoplus_{p\geq 0}\mathfrak{a}_{p}T^{p}$ is finitely generated as an RR–algebra. Let $\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}$ denote its integral closure, and set $\overline{Y}:=\operatorname{Proj}\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}.$ Then, there exists a finite set of divisorial valuations
@@
arising from prime divisors on $\overline{Y}$ lying over $\mathfrak{m}$, such that for every filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in $R$ satisfying $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $1\leq i\leq r$, one has
@@
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of $\mathfrak{m}$-primary ideals whose Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated. Then, there exists a finite set of divisorial valuations $\mathcal{V}$ with the following property:
@@
By the proof of Theorem 5.9, there exists a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})$ computing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ for all filtrations $\mathfrak{b}_{\bullet}$}. The valuative criterion for integral closure then implies that the same set of valuations detects all containments $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$, proving the claim. ∎
@@ 5.5. From supremum to maximum: normalized valuation space and linearly boundedness
By the discussion following [12, Definition 1.3] (see also [35, Proposition 5.9] and the discussion that followed), the normalized space $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ is compact in the topology of pointwise convergence on $R$.
Note that, by [35, Lemma 4.1], $v\mapsto v(\mathfrak{m})$ and $v\mapsto v(\mathfrak{q})$ are continuous on $\operatorname{Val}_{\mathfrak{m}}$. Define
@@
We work on the normalized slice $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})=\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ so that $v(\mathfrak{m})=1$ for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. We use the notation and terminology of Favre–Jonsson monograph [19].
[…]
For $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ set $\beta(v):=\alpha(v\wedge\nu)$. Since $\nu_{i}$ lies on the segment $\operatorname{\mathrm{ord}}_{\mathfrak{m}},\nu]$, the skewness parameterization of segments ([19, Thm. 3.26]) gives
$\alpha(v\wedge\nu_{i})=\min\{\beta(v),A_{i}\}.$
Step 2: simple complete ideals attached to the $\nu_{i}$. For each $i$, let $b_{i}:=b(\nu_{i})$ be the generic multiplicity of the divisorial valuation $\nu_{i}$. Consider the positive atomic measure $\rho_{i}:=b_{i}\nu_{i}$ and define
$J_{i}:=I_{\rho_{i}}=\{\phi\in R\mid g_{\phi}\geq g_{\rho_{i}}\},$
where $I_{\rho}$ is as in [19, (8.3)]. By [19, Thm. 8.12], $J_{i}$ is an integrally closed ideal and its tree measure satisfies
$\rho_{J_{i}}=\rho_{i}.$
Together with [19, Prop. 8.9] and the fact that the maps $I\mapsto\rho_{I}$ and $\rho\mapsto I_{\rho}$ are inverse semigroup isomorphisms between integrally closed ideals and measures in $\mathcal{M}_{\mathcal{I}}^{+}$ ([19, Thm. 8.12]), this also implies that $g_{J_{i}}=g_{\rho_{i}}$ as the corresponding positive tree potential. In particular, by [19, Thm. 7.48 and (8.1)], for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$,
$v(J_{i})=g_{J_{i}}(v)=g_{\rho_{i}}(v)=\int\alpha(\mu\wedge v)\,d\rho_{i}(\mu)=b_{i}\alpha(v\wedge\nu_{i}).$
Furthermore, by the unique factorization of $J_{i}$, as described in [19, Thm. 8.12], and the fact that $\rho_{J_{i}}=b_{i}\nu_{i}$, it follows that this factorization of $J_{i}$ has only one divisorial atom and no curve atoms. Therefore, $J_{i}$ is the simple complete ideal attached to the divisorial valuation $\nu_{i}$ and, in particular, is $\mathfrak{m}$-primary. Particularly, there exists $\ell_{i}\geq 1$ such that
$\mathfrak{m}^{\ell_{i}}\subseteq J_{i}.$
Step 3: integer weights. Choose integers $n_{i}\geq 1$ inductively so that $D_{i}:=\lfloor n_{i}b_{i}A_{i}\rfloor$ is strictly increasing and
$c_{i}:=\frac{n_{i}b_{i}A_{i}}{D_{i}}\longrightarrow 1.$
Set $s_{i}(v):=v(J_{i}^{n_{i}})/D_{i}$. Using the previous formula for $v(J_{i})$,
$s_{i}(v)=\frac{n_{i}b_{i}}{D_{i}}\,\alpha(v\wedge\nu_{i})=\frac{n_{i}b_{i}}{D_{i}}\,\min\{\beta(v),A_{i}\}=c_{i}\frac{\min\{\beta(v),A_{i}\}}{A_{i}}.$
We now prove that $s_{i}(v)\to g(v):=\frac{\beta(v)}{A}=\frac{\alpha(v\wedge\nu)}{A}$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Indeed, for any $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, we have
$\left|s_{i}(v)-\frac{\beta(v)}{A}\right|\leq|c_{i}-1|\frac{\min\{\beta(v),A_{i}\}}{A_{i}}+\left|\frac{\min\{\beta(v),A_{i}\}}{A_{i}}-\frac{\beta(v)}{A}\right|.$
Since $\frac{\min\{\beta(v),A_{i}\}}{A_{i}}\leq 1$ and $0\leq\frac{\min\{\beta(v),A_{i}\}}{A_{i}}-\frac{\beta(v)}{A}\leq 1-\frac{A_{i}}{A},$ we obtain
$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})}\left|s_{i}(v)-g(v)\right|\leq|c_{i}-1|+\left(1-\frac{A_{i}}{A}\right).$
Because $c_{i}\to 1$ and $A_{i}\uparrow A$, the right-hand side tends to 0, so $s_{i}(v)\to g(v)$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$.
Step 4: the auxiliary filtration $\mathfrak{c}_{\bullet}$. For $p\geq 1$ define
$\mathfrak{c}_{p}:=\sum_{\sum_{i}m_{i}D_{i}\geq p}\prod_{i}J_{i}^{n_{i}m_{i}},$
the sum taken over finitely supported tuples $(m_{i})$. It can be seen that $\mathfrak{c}_{\bullet}$ is a filtration of ideals.
Set ${\mathfrak{b}_{\bullet}}:=\mathfrak{m}^{\bullet}$ and $\mathfrak{a}_{p}:=\mathfrak{m}^{p}\mathfrak{c}_{p}$. Then, $\mathfrak{a}_{\bullet}$} is also a filtration of ideals.
Step 5: computation of $v(\mathfrak{c}_{\bullet})$. For any monomial summand $M=\prod_{i}J_{i}^{n_{i}m_{i}}$ with $\sum_{i}m_{i}D_{i}\geq p$,
$v(M)=\sum_{i}m_{i}D_{i}\,s_{i}(v)\geq p\,\inf_{i}s_{i}(v).$
Conversely $J_{i}^{n_{i}\lceil p/D_{i}\rceil}$ occurs among the summands of $\mathfrak{c}_{p}$, so $v(\mathfrak{c}_{p})\leq\lceil p/D_{i}\rceil v(J_{i}^{n_{i}})$. We shall show that
(5.4) $s_{i}(v)\geq g(v)\text{ for all }i\text{ and all }v.$
Indeed, observe that $s_{i}(v)\geq\frac{\min\{\beta(v),A_{i}\}}{A_{i}}.$ If $\beta(v)\leq A_{i}$, then
$\frac{\min\{\beta(v),A_{i}\}}{A_{i}}=\frac{\beta(v)}{A_{i}}\geq\frac{\beta(v)}{A}=g(v),$
whereas if $\beta(v)\geq A_{i}$, then
$\frac{\min\{\beta(v),A_{i}\}}{A_{i}}=1\geq\frac{\beta(v)}{A}=g(v).$
This proves (5.4).
Next, as shown above, $s_{i}(v)\longrightarrow g(v)$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Therefore, together with (5.4), we have
$\inf_{i}s_{i}(v)=g(v)\text{ for all }v.$
Now, it can be seen that $v(\mathfrak{c}_{p})\geq pg(v)$. On the other hand, for each fixed $i$, $J_{i}^{n_{i}\lceil p/D_{i}\rceil}\subseteq\mathfrak{c}_{p},$ so $v(\mathfrak{c}_{p})\leq\left\lceil\frac{p}{D_{i}}\right\rceil v(J_{i}^{n_{i}}),$ and therefore,
$\frac{v(\mathfrak{c}_{p})}{p}\leq s_{i}(v)+\frac{n_{i}b_{i}A_{i}}{p}.$
We shall prove
$\lim_{p\to\infty}\frac{v(\mathfrak{c}_{p})}{p}=g(v)$
uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Let $\varepsilon>0$. By the uniform convergence of $s_{i}$ to $g$ proved before, choose $i_{0}$ such that
$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})}|s_{i_{0}}(v)-g(v)|<\varepsilon/2.$
Since $v(J_{i_{0}}^{n_{i_{0}}})\leq n_{i_{0}}b_{i_{0}}A_{i_{0}}$ for all $v$, choose $p_{0}$ such that
$\frac{n_{i_{0}}b_{i_{0}}A_{i_{0}}}{p}<\varepsilon/2\text{ for all }p\geq p_{0}.$
For such $p$, the lower bound above gives
$\frac{v(\mathfrak{c}_{p})}{p}\geq g(v),$
while the upper bound with $i=i_{0}$ gives
$\frac{v(\mathfrak{c}_{p})}{p}\leq s_{i_{0}}(v)+\frac{n_{i_{0}}b_{i_{0}}A_{i_{0}}}{p}\leq g(v)+\varepsilon$
for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Thus,
$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})}\left|\frac{v(\mathfrak{c}_{p})}{p}-g(v)\right|\leq\varepsilon$
for all $p\geq p_{0}$, establishing the desired uniform convergence.
Step 6: valuation of $\mathfrak{a}_{\bullet}$}. Since $\mathfrak{a}_{p}=\mathfrak{m}^{p}\mathfrak{c}_{p}$ and $v(\mathfrak{m})=1$,
$\frac{v(\mathfrak{a}_{p})}{p}\to 1+g(v)=1+\frac{\alpha(v\wedge\nu)}{A}.$
Hence, $v({\mathfrak{a}_{\bullet}})=1+\alpha(v\wedge\nu)/A$ and $v({\mathfrak{b}_{\bullet}})=1$.
Step 7: maximizer. At $v=\nu$ this ratio equals $2$. If $v\neq\nu$, then $v\wedge\nu\prec\nu$ and by strict monotonicity of skewness along segments ([19, Thm. 3.26]), $\alpha(v\wedge\nu)<A$. It follows that $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})<2=\nu({\mathfrak{a}_{\bullet}})/\nu({\mathfrak{b}_{\bullet}})$. Thus, $\nu$ uniquely maximizes $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})$, and since $\nu$ is not divisorial, no divisorial valuation attains the supremum.
@@
Step 8: linear boundedness. Because $\mathfrak{c}_{p}\subseteq R$, one has $\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}$. Also $J_{1}^{n_{1}\lceil p/D_{1}\rceil}\subseteq\mathfrak{c}_{p}$ and $\mathfrak{m}^{\ell_{1}}\subseteq J_{1}$, so $\mathfrak{m}^{p+n_{1}\ell_{1}\lceil p/D_{1}\rceil}\subseteq\mathfrak{a}_{p}$. Thus $\mathfrak{m}^{\lceil cp\rceil}\subseteq\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}$ for some $c>0$, and $\mathfrak{a}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$. ∎